# Governing Together: Toward Infrastructure for Community-Run Social Media

### v1

@@ 4.2. Analytical Approach
Figure 2. From the Miro boards, a synthesized summary of community decisions listed from the four workshops.:
@@
The first author began by reviewing the boards, [^3] A screenshot of one board is shown in the Appendix as an example. synthesizing across the sessions. We compiled a list of community decisions and ideas for social, organizational, and technical tools that participants discussed. Community decisions, shown in Figure 2, were clustered into groupings that might reflect broader dimensions of community governance. Tools were grouped based on similarity of ideas (e.g., some ideas were recurrent or highly similar). These syntheses provided a rough map of topics and ideas across all participants, discussed in a team-wide meeting in July 2025.
We followed Braun and Clarke (2006)’s thematic analysis approach to identify challenges in inter-community governance. The first author conducted inductive line-by-line coding of the workshop transcripts, with an eye towards identifying motivations and concerns the tools responded to as well as surfaced. Moving between the transcripts and the synthesized mapping of ideas, the first author iteratively grouped the tools into core challenges. Four authors directly involved with data collection and cleaning met to discuss the stability of these groupings, agreeing on five major themes that felt both substantial and distinct. Then, the first author inductively coded the follow-up interviews, with the intent of stress-testing the set of themes. This resulted in the development of one final theme that had not been captured previously.
@@ 5. Findings
We organize our findings around six distinct challenges for inter-community governance. The first three sections focus on issues around information-seeking; the next two, on how relationships between communities are subsequently managed; and the final section, on questions around how interventions can be effective.
@@ 5.1. Making Decisions About Community Governance Visible
Participants across workshops noted that governance decisions (e.g., see Figure 2) by communities were often invisible to people outside of the community’s admin and mod team, making it difficult to anticipate differences — and any consequences. They thus frequently proposed mechanisms that might make a community’s governance decisions more salient. For example, in Workshop 4, P17 proposed a community governance nutrition facts label, inspired by existing tools for selecting open source licenses. [^4] https://choosealicense.com/ P17 imagined such a label might describe decisions a community made along distinct dimensions of community governance — its moderation process, a link to its rules, its decision-making model (e.g., coop, single admin, etc.), its goals for connectivity, and so on. This would make it easier for prospective members to assess a community, providing “much stronger signals” (P5) by enabling direct comparisons.
Participants suggested that consistent categories provide a “shared language/vocabulary” (P5, P10, P11, P12, P14) that offer explanation about a community’s behavior toward reducing tensions about differences. A lack of context about the basic goals of a community, or the size of its admin team, could lead to questions about “why the content is moderated the way it is […] and why they’re seeing it in that space,” producing “a lot of divide” (P17).
Making dimensions of governance salient to others could also become a guide for running one’s own community. As P9 remarked, this was crucial because each community impacts others: “you’re not setting up just your little corner of the web. You are actually joining this - you are proposing to make a node on this network. […] Think about why your community exists. How can you reflect your community in these rules?” Participants speculated that surfacing dimensions of governance could prompt intentional community decisions, such as around federation or moderation processes. P21 provided an analogy:
@@ 5.2. Collating Nuanced Information About Issues
Participants wanted to know what issues other communities were dealing with, or had dealt with in the past. This was discussed in context of ideas such as a moderator bulletin or shared modbox (moderator inbox), envisioned to look across many communities to “pull together [public] information and make it more digestible” (P1) so that moderators could “handle the flood of crap” (P20).
@@
As participants noted, some conceptually-similar tools and resources already exist. Participants such as P1, P10, and P16 referenced IFTAS Connect, a forum for moderators maintained by a non-profit (IFTAS, or Independent Federated Trust and Safety). [^5] https://about.iftas.org/ Other participants pointed to more private groups of moderators using Discord, email listservs, and other communication channels. A widespread existing resource mentioned across workshops was the #Fediblock hashtag, which had been started by a group of queer femmes (Kiam, 2023) to help moderators across communities alert one another about bad actors. Although these resources were described as valuable, the information they provided was usually not systematic or contextualized enough unless one was “following this stuff extremely closely” (P1). For example, an individual might post about a community they had a bad experience with on the #Fediblock hashtag, but not include receipts or why they felt the community should be blocked.
@@
Participants thus wanted bulletins and inboxes to automatically present consistent, structured, and nuanced information about issues: which ones were important, what were their main facts, how others were approaching them. Participants proposed these tools would pool information about incidents, possibly through automatically shared moderation logs that were processed into discrete issues. Issues then might be algorithmically presented in order of “trending” problems. Participants imagined being able to get deeper context about specific cases, reports, and content, sometimes through corresponding discussion threads where they could “sort of share and bounce ideas and like, how would you approach this?’’ (P10). P4 described what they imagined:
@@
Here’s the hot issues coming up. For this issue, here’s the receipts that I’ve seen, here’s the context that I’ve seen. You know? Then people add in — “Oh no, there’s this other important discussion here. Oh, here’s how person A is thinking about it”. So … A loosely structured, almost a separate timeline with more pooled information about: What’s important? What should I know about it? And what are other people thinking about it in my circle?
@@
Such a tool was imagined to help make well-grounded, thoughtful decisions efficiently and accurately. For example, participants described wanting to distinguish between harm by an individual vs a whole community: ‘Ideally, situations like these can be handled, like, pretty surgically, where whoever the individual bad actors are, they [individually] can just get removed” (P2). Such an example emphasizes being able to find enough context about a case to make fair, nuanced judgments of who to hold accountable when harm occurs.
@@ 5.3. Managing What Information is Shared Across Trust Bubbles
Participants brainstormed how information might be shared in tiered ways through concepts such as trust bubbles: smaller networks or coalitions of other communities that they felt like they could trust. In sketching out the bulletins, modboxes, and other imagined interfaces for information-sharing, participants emphasized that these spaces did not need to contain everyone in the network. A community might also have multiple trust bubbles that they moved across, with some having more trust — and more access to information — than others. P6 put it succinctly: “I would really like that an admin [has] a trusted coalition that they can work with in total trust, or across different levels of trust — because you don’t have to share everything and all the details, but you can have different levels of trust where you talk, discuss […] to try to address issues from a multifaceted point of view.
@@
Discussions touched on how sharing with broader or less trusted bubbles of communities would work. Participants proposed identifying abstract signals could still give people information about problems and other communities. P3 suggested “some automated tool where maybe [it checks if] conversations have a different structure” that might indicate reason for concern, examining “the shape of data and the shape of conversations”, but not retaining their content. P16 spit-balled on more straightforward metrics, such as time to resolution for reports. Signals needed to give individuals a sense of the relevance of a potential problem to their community and the urgency with which they should act. They also wanted to be able to share signals of their own decisions or ongoing efforts to prevent misunderstandings or unnecessary conflict. P4, for example, explained that because their community valued democratic decision-making, they were often slower in dealing with cases: “Our [server] could get blocked just because we have a longer process […] And it would be great if we could communicate that [we’re dealing with it] to the whoever it was that reported the issue.” While one might be able to just be fully transparent with others in very close trust bubbles, abstract signals could both protect the privacy of community decisions and help communicate about them.
@@ 5.4. Enabling New Relationships Between Communities
To incentivize communities toward cooperation, participants suggested social mechanisms that might make communities feel accountable to one another. As noted in §4.1, each workshop had opened with a discussion about cooperation among fishers to prevent overfishing of lobsters in Maine. Participants pointed out how the shared economic interests of the fishers, compared to the diverse interests on the Fediverse, drew them to be responsible to one another. In turn, participants considered how relationships among a set of communities — i.e., a trust bubble — might be configured to produce a similar effect: “ [The bubble] agree[s] to a minimal amount of shared rules and enforcement. And if you don’t, then you might get removed from the [bubble’s] list. […] But then you could have alliances between alliances too, making it a true federation where it’s not just either you’re in or you’re out” (P15).
The quote above suggests moving beyond a dichotomy of “either you’re in or you’re out” in how communities engage with each other. This was further fleshed out in the idea of ladder of engagement, raised by P16: “I start with a certain amount of trust — or, like, start silenced — and […] if I want to federate with another instance, I would have to be like, okay, cool, I get this amount of federation by default, and then if I can prove to them [that I’m trustworthy], then I can get upgraded to platinum status.” The ladder of engagement might be a conceptual — and encoded — means through which the trust bubbles that participants discussed are formed, sustained, become tiered, and evolve.
Participants hoped that such mechanisms would enable more nuanced types of relationships between communities beyond the federation-defederation dichotomy. P13 said:
@@ 5.5. Customizing Community Governance Decisions
As participants valued community autonomy, the shared nature of tools and resources often raised concerns that the Fediverse would perceive them as centralizing or homogenizing forces. For example, the community governance nutrition fact labels (§5.1) might be seen as demanding communities to conform. P9 humorously likened it to being a homeowner’s association: “Lots of servers have rules that are like, Everybody, be nice! Golden Rule. That’s all you need. And, okay — how do we perhaps suggest that you want a little bit more of [substance in] that, without sounding like we’re the homeowners association coming in and saying: Nope, nope, your rules are all bad, you need to adopt ours instead?
Participants thus emphasized the importance of first developing a consistent, shared language about community governance that could prompt communities to make well-considered, independent decisions. A few participants imagined this as being implemented through a forkable, modular governance system — perhaps underpinning the generation of community governance nutrition facts — where a community could search through a “user-friendly […] set of options” (P12) containing decisions, practices, and processes that could be copied and tweaked from other communities. Participants envisioned that this system would cover different dimensions of governance (“How do you structure a moderation team? How do you organize? What’s the process when someone flags a post?” (P15), encoding them to be human and machine-readable. As communities went through the dimensions, P12 imagined that they would be prompted to reflect how a choice reflected its unique goals and needs:
@@
5.6. Minimizing Technical Barriers to Inter-Community Governance
@@
Participants suggested that the best way to overcome this was to make any technical or infrastructural tools easily integrated or built-in to existing technologies so that people could avoid “splitting [their] energy and focus” (P6). Participants such as P4, P13, and P23 noted that they already had too many tools that their community was juggling. This made logging into and learning to use yet another tool for inter-community governance unappealing, especially as the need for inter-community governance is not always obvious in the daily work of moderators. At the same time, when questions or harms arise, the benefits of and need for inter-community governance become clear. One intuitive path forward might be to foster relationships with developers of Fediverse platforms used to run a server, to encourage them to build the kinds of tools participants described in workshops since such software “determines a lot of [what] different governance models become available to you” (P11).
@@
Having been encouraged to think about organizational and social approaches, participants also questioned whether a “tool” had to be technical in nature. At the core, any tools devised to support inter-community governance had to be “social tools” (P5). P4 stated: “I think that really, at a deeper level, this is about moderators across communities talking to each other and building enough of a community that we’re able to really be in touch and know who to talk to and when to have these conversations.” Participants pointed out that simply having regular open moderator calls might help foster the community of communities they envisioned. P13 imagined how this simple sense of community might in turn help encourage the adoption of technical tools that made governance work easier: “Wouldn’t it be nice if there was a tool […] and I announced [it] to my trust bubble: Hey, did you see this thing? […] And now, it bops, bops, bops [throughout the trust bubbles], and suddenly 50% of the Fediverse is doing it, and the rest can do it by osmosis.” In short, a broader cultural shift might be facilitated by technologies — but ultimately, does not require them.
@@ 6. Discussion
Figure 3. Here we visualize the approaches imagined by participants as an ecosystem of shared resources, spaces, and automated tooling. A community (C) shows its “Governance Nutrition Facts” label, which can be generated through a user-friendly interface that modularizes dimensions of community governance. Communities can fork and customize these labels. Communities sit in polycentric “trust bubbles,” or coalitions of communities, that are indicated by the colored clouds grouping communities. These trust bubbles help structure how communities share and receive information about governance from one another, shown in an “Issue Bulletin” for each trust bubble. The Issue Bulletin automatically processes governance activity logs into a database a community can review. It also enables communities to engage in discussion.:

### v2

@@ 4.2. Analytical Approach
The first author began by reviewing the boards, [^3] A screenshot of one board is shown in the Appendix as an example. synthesizing across the sessions. We compiled a list of community decisions and ideas for social, organizational, and technical tools that participants discussed. Community decisions raised in the workshops, shown in Table 2 in Appendix B, were clustered into groupings that might reflect broader dimensions of community governance. Tools were grouped based on similarity of ideas (e.g., some ideas were recurrent or highly similar) and are shown in Table 3 in Appendix B. These syntheses provided a rough map of topics and ideas across all participants, discussed in a team-wide meeting in July 2025.
We followed Braun and Clarke (2006)’s thematic analysis approach to identify challenges in inter-community governance. The first author conducted inductive line-by-line coding of the workshop transcripts, with an eye towards identifying motivations and concerns the tools responded to as well as surfaced. Moving between the transcripts and the synthesized mapping of ideas, the first author iteratively grouped the tools into core challenges. Four authors directly involved with data collection and cleaning met to discuss the stability of these groupings, agreeing on five major themes that felt both substantial and distinct. Then, the first author inductively coded the follow-up interviews, stress-testing the set of themes. This resulted in the development of one final theme that had not been captured previously.
@@ 5. Findings
We organize our findings around six distinct challenges for inter-community governance. The first three sections focus on issues around information-seeking about and from other communities; the next two, on how relationships between communities are subsequently managed; and the final section, on questions around how interventions can be effective.
@@ 5.1. Making Decisions About Community Governance Visible
Participants across workshops noted that governance decisions (e.g., Table 2) by communities were often invisible to people outside of the community’s admin and mod team, making it difficult to anticipate differences — and any frictions that arose. They thus frequently proposed mechanisms that might make a community’s governance decisions more salient. For example, in Workshop 4, P17 proposed a community governance nutrition facts label, inspired by existing tools for selecting open source licenses. [^4] https://choosealicense.com/ P17 imagined such a label might describe decisions a community made along distinct dimensions of community governance — its moderation process, a link to its rules, its decision-making model (e.g., coop, single admin, etc.), its goals for connectivity, and so on. Consistent structure would make it easier for people to assess a community, providing “much stronger signals” (P5) by enabling direct (and possibly automated) comparisons across a network of communities.
Participants suggested that consistent categories provide a “shared language/vocabulary” (P5, P10, P11, P12, P14) that offer explanation about a community’s behavior toward reducing tensions about frictions. A lack of context about the basic goals of a community, or the size of its admin team, could lead to questions about “why the content is moderated the way it is […] and why they’re seeing it in that space,” producing “a lot of divide” (P17).
Making dimensions of governance salient to others could also become a guide for running one’s own community. As P9 remarked, this was crucial because each community impacts others: “you’re not setting up just your little corner of the web. You are actually joining this - you are proposing to make a node on this network. […] Think about why your community exists. How can you reflect your community in these rules?” Participants speculated that surfacing dimensions of governance could prompt intentional community decisions, such as around federation or moderation. P21 provided an analogy:
@@ 5.2. Collating Nuanced Information About Issues
Across workshops, participants wanted to know what issues other communities were dealing with, or had dealt with in the past, to anticipate real-time problems and get advice from others. This was often discussed in context of ideas such as a moderator bulletin or shared modbox (moderator inbox), envisioned to look across many communities to “pull together [public] information and make it more digestible” (P1) so that moderators could “handle the flood of crap” (P20).
@@
As participants noted, some conceptually-similar spaces and resources already exist. Participants such as P1, P10, and P16 referenced IFTAS Connect, a forum for moderators maintained by a non-profit (IFTAS, or Independent Federated Trust and Safety). [^5] https://about.iftas.org/ Other participants pointed to more private groups of moderators using Discord, email listservs, and other communication channels. A widespread existing resource mentioned across workshops was the #Fediblock hashtag, which had been started by a group of queer femmes (Kiam, 2023) to help moderators across communities alert one another about bad actors. Although these resources were described as valuable, the information they provided was usually not systematic or contextualized enough unless one was “following this stuff extremely closely” (P1). For example, an individual might post about a community they had a bad experience with on the #Fediblock hashtag, but not include evidence or an explanation.
@@
Participants thus wanted bulletins and inboxes to automatically present relevant, structured, and nuanced information about issues: which ones were important, what were their main facts, how others were approaching them. Participants proposed pooling information about incidents, possibly through automatically shared moderation logs that were processed into discrete issues. Issues then might be algorithmically presented in order of “trending” problems or relevance for their community. Participants imagined being able to get deeper context about specific cases, reports, and content, sometimes through corresponding discussion threads where they could “sort of share and bounce ideas and like, how would you approach this?’’ (P10). P4 brainstormed:
@@
Here’s the hot issues coming up. For [an] issue, here’s the receipts that I’ve seen, here’s the context that I’ve seen. You know? Then people add in — “Oh no, there’s this other important discussion here. Oh, here’s how person A is thinking about it”. So … A loosely structured, almost a separate timeline with more pooled information about: What’s important? What should I know about it? And what are other people thinking about it in my circle?
@@
Such a tool was imagined to help proactively anticipate problems in the network, as well as help community leaders learn from another. Participants also imagined having information through these spaces could help them make well-grounded decisions. For example, participants described wanting to distinguish between harm done by an individual vs a whole community: “Ideally, situations like these can be handled, like, pretty surgically, where whoever the individual bad actors are, they [individually] can just get removed” (P2). Being able to find enough context about a case would enable communities to make fair, nuanced judgments of who to hold accountable. For example, rather than banning an entire community, one might message its leaders to note that some of their rules are too broad and have allowed bad actors to slip by.
@@ 5.3. Managing What Information is Shared Across Trust Bubbles
In all workshops, participants brainstormed how information might be shared in tiered ways through concepts, most explicitly named in Workshop 3 as trust bubbles: smaller networks or coalitions of other communities that they felt like they could trust. In sketching out the bulletins, modboxes, and other imagined interfaces for information-sharing, participants emphasized that these spaces did not need to contain everyone in the network. A community might also have multiple trust bubbles that they moved across, with some having more trust — and more access to information — than others depending on the nature of governance frictions they had. P6 put it succinctly: “I would really like that an admin [has] a trusted coalition that they can work with in total trust, or across different levels of trust — because you don’t have to share everything and all the details, but you can have different levels of trust where you talk, discuss […] to try to address issues from a multifaceted point of view.
@@
Discussions in Workshops 1 and 4 in particular touched on how sharing with more vs less trusted bubbles of communities would work, possibly through abstract signals that could still give people information about problems and other communities. P3 suggested “some automated tool where maybe [it checks if] conversations have a different structure” that might indicate reason for concern, examining “the shape of data and the shape of conversations”, but not retaining their content. P16 spit-balled on more straightforward metrics that could be proxies, such as time to resolution for reports. Signals needed to give individuals a sense of the relevance of a potential problem and the urgency with which they should act. They also wanted to be able to share signals of their own decisions or ongoing efforts to prevent misunderstandings or unnecessary conflict if frictions emerged. P4, for example, explained that because their community valued democratic decision-making, they were often slower in dealing with cases: “Our [server] could get blocked just because we have a longer process […] And it would be great if we could communicate that [we’re dealing with it] to the whoever it was that reported the issue.” While one might be able to just be fully transparent with others in very close trust bubbles, abstract signals could both protect the privacy of community decisions and help communicate about them.
@@ 5.4. Enabling New Relationships Between Communities
To incentivize communities toward cooperation, participants suggested social mechanisms that might make communities feel accountable to one another, possibly inspired by the opening example. As noted in §4.1, each workshop had opened with a discussion about cooperation among fishers to prevent overfishing of lobsters in Maine. Participants pointed out how the shared economic interests of the fishers who lived and worked in the same region, compared to the diverse interests on the Fediverse, drew them to be responsible to one another. In turn, participants considered how relationships among a set of communities — i.e., a trust bubble — might be configured to produce a similar effect: “ [The bubble] agree[s] to a minimal amount of shared rules and enforcement. And if you don’t, then you might get removed from the [bubble’s] list. […] But then you could have alliances between alliances too, making it a true federation where it’s not just either you’re in or you’re out” (P15).
The quote above suggests moving beyond a dichotomy of “either you’re in or you’re out” in how communities engage with each other. This was particularly fleshed out in the idea of ladder of engagement, raised by P16 in Workshop 4: “I start with a certain amount of trust — or, like, start silenced — and […] if I want to federate with another instance, I would have to be like, okay, cool, I get this amount of federation by default, and then if I can prove to them [that I’m trustworthy], then I can get upgraded to platinum status.” The ladder of engagement might be a conceptual — and encoded — means through which the trust bubbles that participants discussed are formed, sustained, become tiered, and evolve.
Participants hoped that such mechanisms would enable more nuanced types of relationships between communities beyond a federation-defederation dichotomy. P13 said:
@@ 5.5. Customizing Community Governance Decisions
As participants valued community autonomy, the shared nature of tools and resources raised concerns that the Fediverse would perceive them as centralizing or homogenizing forces. For example, the community governance nutrition fact labels (§5.1) might be seen as demanding communities to conform. P9 humorously likened it to being a homeowner’s association: “Lots of servers have rules that are like, Everybody, be nice! Golden Rule. That’s all you need. And, okay — how do we perhaps suggest that you want a little bit more of [substance in] that, without sounding like we’re the homeowners association coming in and saying: Nope, nope, your rules are all bad, you need to adopt ours instead?
Participants thus emphasized the importance of first developing a consistent, shared language about community governance that could prompt communities to make well-considered, independent decisions. A few participants in Workshop 3 imagined this as being implemented through a forkable, modular governance system — perhaps underpinning the generation of community governance nutrition facts — where a community could search through a “user-friendly […] set of options” (P12) containing decisions, practices, and processes that could be copied and tweaked from other communities. Participants envisioned that this system would cover different dimensions of governance (“How do you structure a moderation team? How do you organize? What’s the process when someone flags a post?” (P15), encoding them to be human and machine-readable. As communities went through the dimensions, P12 imagined that they would be prompted to reflect how a choice reflected its unique goals and needs:
@@
5.6. Minimizing Barriers to Inter-Community Governance
@@
Participants suggested that the best way to overcome this was to make any technical or infrastructural tools easily integrated or built-in to existing technologies so that people could avoid “splitting [their] energy and focus” (P6). Across workshops, participants such as P4, P13, and P23 noted that they already had too many tools that their community was juggling. This made logging into and learning to use yet another tool for inter-community governance unappealing, especially as the need for inter-community governance is not always obvious in the daily work of moderators. At the same time, when questions or harms arise, the benefits of and need for inter-community governance become clear. One intuitive path forward might be to foster relationships with developers of Fediverse platforms used to run a server, to encourage them to build the kinds of tools participants described in workshops since such software “determines a lot of [what] different governance models become available to you” (P11).
@@
Having been told that a tool could include organizational and social approaches, participants also questioned whether a “tool” had to be technical in nature. At the core, any tools devised to support inter-community governance had to be “social tools” (P5). P4 stated: “I think that really, at a deeper level, this is about moderators across communities talking to each other and building enough of a community that we’re able to really be in touch and know who to talk to and when to have these conversations [about frictions].” Participants pointed out that simply having regular open moderator calls might help foster the community of communities they envisioned. P13 imagined how this simple sense of community might in turn help encourage the adoption of technical tools that made dealing with frictions easier: “Wouldn’t it be nice if there was a tool […] and I announced [it] to my trust bubble: Hey, did you see this thing? […] And now, it bops, bops, bops [throughout the trust bubbles], and suddenly 50% of the Fediverse is doing it, and the rest can do it by osmosis.” In short, a cultural shift might be facilitated by technologies — but ultimately, does not require them.
@@ 6. Discussion
Figure 3. Visualization of the approaches imagined by participants as an ecosystem of shared resources, spaces, and automated tooling. A community (C) shows its “Governance Nutrition Facts” label, which can be generated through a user-friendly interface that modularizes dimensions of community governance. Communities can fork and customize these labels. Communities sit in polycentric “trust bubbles,” or coalitions of communities, that are indicated by the colored clouds grouping communities. These trust bubbles help structure how communities share and receive information about governance from one another, shown in an “Issue Bulletin” for each trust bubble. The Issue Bulletin automatically processes governance activity logs into a database a community can review. It also enables communities to engage in discussion.
The figure gives a conceptual overview of all the different ideas proposed by participants, in one ecosystem. It shows ten gray circles, each representing a community. Three colored blobs covering different but overlapping subsets of them indicate the ‘trust bubble’ they are in. The figure shows how communities each have Governance Nutrition Facts, which looks like a label with a consistent icons on the left representing a dimension of community governance. Presumably, communities note their decision for each dimension. On the right of the labels, there is a note about modular forking, which enables communities to copy and adapt the Governance Nutrition Facts from one another. The image also shows a community exporting its activity log to an ‘Issue Bulletin,’ a shared space for the community’s trust bubble. The Bulletin has a split pane, where the left pane shows aggregated issues and meta-information about it and the right pane shows discussion threads. Notably, left of the Issue Bulletin are three dots that correspond to the trust bubble blobs. Communities can select which trust bubble’s Issue Bulletin they want to see and contribute to. The figure also has a note about the ‘ladder of engagement,’ which is a concept to ground how communities change their information sharing depending on where on the ladder they are with others. For example, communities that are mutually low on the ladder of engagement may only share abstract metadata. Communities that are mutually high on the ladder of engagement might send much more detailed information about their processes.